OpenStudy (anonymous):

Solve. 19 + 8c = −5 − 33c

OpenStudy (blank ):

c = -24/41

OpenStudy (blank ):

Tell me if you need the steps.

OpenStudy (anonymous):

yes plz

OpenStudy (blank ):

19+8c=-5-33c Move all terms not containing c to the right-hand side of the equation. 41c+19=-5 Move all terms not containing c to the right-hand side of the equation. 41c=-24 Divide each term in the equation by 41. c=-(24)/(41) Hope this helps. :)

OpenStudy (anonymous):

ty one more BAn amusement park has two types of season passes. Plan 1 charges a one-time fee of $175.00 for admission plus $8.00 every trip for parking. Plan 2 charges a one-time fee of $125.00 for parking plus $13.00 every trip for admission. For what number of trips is the cost of these plans the same?

OpenStudy (blank ):

Just a second...

OpenStudy (anonymous):

k

OpenStudy (blank ):

10 trips

OpenStudy (anonymous):

ty ty

OpenStudy (anonymous):

An amusement park has two types of season passes. Plan 1 charges a one-time fee of $170.00 for admission plus $8.00 every trip for parking. Plan 2 charges a one-time fee of $93.00 for parking plus $15.00 every trip for admission. For what number of trips is the cost of these plans the same?

OpenStudy (blank ):

What are your choices?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

5 trips 11 trips 9 trips 21 trips

OpenStudy (blank ):

11 trips

OpenStudy (anonymous):

ty lat one i have to do An amusement park has two types of season passes. Plan 1 charges a one-time fee of $145.00 for admission plus $5.00 every trip for parking. Plan 2 charges a one-time fee of $40.00 for parking plus $12.00 every trip for admission. For what number of trips is the cost of these plans the same? 8 trips 15 trips 21 trips 29 trips

OpenStudy (blank ):

15 trips

OpenStudy (anonymous):

im A FAN you good at history

OpenStudy (blank ):

No, not really. Sorry. :)